# ReaCritic: Large Reasoning Transformer-based DRL Critic-model Scaling For Heterogeneous Networks

### v1

@@ Abstract
Heterogeneous Networks (HetNets) pose critical challenges for intelligent management due to the diverse user requirements and time-varying wireless conditions. These factors introduce significant decision complexity, which limits the adaptability of existing Deep Reinforcement Learning (DRL) methods. In many DRL algorithms, especially those involving value-based or actor-critic structures, the critic component plays a key role in guiding policy learning by estimating value functions. However, conventional critic models often use shallow architectures that map observations directly to scalar estimates, limiting their ability to handle multi-task complexity. In contrast, recent progress in inference-time scaling of Large Language Models (LLMs) has shown that generating intermediate reasoning steps can significantly improve decision quality. Motivated by this, we propose ReaCritic, a large reasoning transformer-based critic-model scaling scheme that brings reasoning ability into DRL. ReaCritic performs horizontal reasoning over parallel state-action inputs and vertical reasoning through deep transformer stacks. It is compatible with a broad range of value-based and actor-critic DRL algorithms and enhances generalization in dynamic wireless environments. Extensive experiments demonstrate that ReaCritic improves convergence speed and final performance across various HetNet settings and standard OpenAI Gym control tasks.
@@
I Introduction
The rapid growth of intelligent wireless services such as the Internet of Things (IoT), Mobile Edge Computing (MEC), and smart city applications is driving the evolution of future wireless networks toward large-scale Heterogeneous Networks (HetNets) [1]. These HetNets involve a large number of mobile users and dynamic environments with heterogeneous communication and computing resources. As a result, HetNets must handle massive data traffic, ensure low-latency responses, and maintain energy efficiency under diverse service demands, which introduces fundamental challenges for resource management and decision-making [2]. In particular, the coexistence of heterogeneous user types, diverse Quality-of-Service (QoS) requirements, and time-varying wireless conditions leads to high-dimensional state spaces, complex user-resource interactions, and non-stationary dynamics. These properties hinder the scalability and generalization of traditional optimization and supervised learning-based methods [3].
@@ I Introduction
Faced with these challenges, Deep Reinforcement Learning (DRL) has gained attention as a data-driven approach for autonomous network management in complex wireless environments. By interacting with the environment and optimizing long-term rewards, DRL enables agents to adapt to dynamic network conditions without relying on predefined models [4]. Its success in domains such as intelligent control, robotics, and wireless resource management demonstrates its potential for handling large-scale, uncertain systems [5, 6, 7]. However, when applied to large-scale HetNets, DRL still faces key limitations not only due to environmental complexity but also due to internal architectural bottlenecks. Although DRL algorithms are capable of learning adaptive behaviors through trial-and-error interaction, many existing designs rely on simple Multi-Layer Perceptron (MLP)-based model structures that treat input states as flat feature vectors and apply direct approximations for policy or value functions. This formulation fundamentally lacks the capacity to capture latent dependencies among state variables or to represent tradeoffs inherent in multi-objective decision-making. In heterogeneous wireless environments, where metrics such as latency, throughput, and energy efficiency are often tightly coupled and evolve asynchronously, such limitations can lead to unstable training, slow convergence, and poor performance in dynamic or unseen scenarios. Taking actor-critic-based DRL algorithms as an example, as shown in Fig. 1, the critic network estimates Q-values based on environmental states and the actions selected by the actor network. However, due to the high dimensionality and structural complexity of both the HetNet states and the action space, this mapping becomes difficult and inaccurate, limiting the effectiveness of actor updates and ultimately degrading overall DRL performance.
@@
Therefore, it is essential to enhance the scalability of DRL frameworks to better accommodate high-dimensional state spaces and dynamic network behaviors. One common strategy is to increase model depth or width, with the goal of improving the DRL’s capacity to interpret complex input features. However, scaling up the MLP parameter size alone often leads to overfitting, unstable convergence, and limited robustness in non-stationary environments. A more effective direction is to improve the DRL model through architectural adjustments that enable it to process structured observations and capture latent dependencies within the environment, beyond what can be achieved through conventional model size expansion. In this context, the architectural principles of Large Language Models (LLMs) offer a compelling source of inspiration. Recent advances in reasoning-capable LLMs have demonstrated strong capabilities in pattern abstraction and generalization across a wide range of tasks [8]. These LLMs are particularly effective at modeling intricate relationships and long-term dependencies in structured or semi-structured data, opening new opportunities for improving representation learning and decision-making in complex environments.
@@
  • Q1 (How): How can the reasoning capability of LLMs be effectively adapted to DRL settings, if not directly incorporating pre-trained LLMs as external modules?
@@
To address Q1 (How), we adopt a Transformer-based reasoning architecture as a more stable, scalable, and task-adaptive alternative to conventional MLP-based designs. By leveraging self-attention mechanisms and stacked abstraction layers, the Transformer enables structured reasoning over state-action representations. This allows for more accurate value estimation under multi-objective and dynamic conditions, while avoiding the overfitting and instability often associated with naive MLP scaling. The hierarchical and modular design of the Transformer also facilitates its integration into DRL training pipelines, where robustness and adaptability are essential. To address Q2 (Where), we embed this reasoning capability into the critic module of the DRL framework. As the critic is responsible for value estimation and provides gradient signals to guide policy updates, it is the natural component for incorporating structured inference. This choice preserves the original policy structure while substantially improving the agent’s ability to model long-term dependencies, reason over complex state-action trajectories, and adapt to heterogeneous and time-varying network environments. Our proposed design, referred to as ReaCritic, is illustrated in Part C of Fig. 1.
@@
  • We propose ReaCritic, a large reasoning transformer-based critic-model scaling architecture for DRL, inspired by inference-time scaling in LLMs. ReaCritic enables horizontal reasoning across parallel state-action token sequences and vertical reasoning through stacked transformer blocks, supporting expressive and adaptable Q value estimation in complex environments.
  • We conduct extensive experiments in a large-scale multi-user HetNet environment to evaluate ReaCritic. Compared with conventional MLP-based critics and standard DRL baselines, ReaCritic achieves faster convergence, improved policy stability, and stronger adaptability under dynamic and heterogeneous network conditions.
The remainder of this paper is organized as follows: Section II discusses related work in HetNets, DRL for networks, and transformer for wireless networks. Section III presents the wireless network system model. Section III-B gives the problem formulation of the optimization problem in the considered HetNet. Section IV details the design of ReaCritic and the reasoning transformer-based critic-model scaling. Section V presents comprehensive experimental results and analysis. Finally, Section VI concludes the paper with a summary of our key findings. A list of mathematical symbols frequently used in this paper is shown in Table I.
@@ II-A Heterogeneous Networks
HetNets have been studied using a range of techniques, including optimization [17], game theory [18], heuristics [19], and more recently, DRL [20]. These approaches aim to address the diverse user demands and dynamic resource constraints inherent in HetNets. Despite progress, the high dimensionality and coupling among decision variables make HetNets a particularly challenging yet critical domain for intelligent network management [21]. In [22], the authors proposed a secrecy-aware, energy-efficient scheme for HetNets, jointly optimizing power control, channel allocation, and beamforming under secrecy and Signal-to-Interference-plus-Noise Ratio (SINR) constraints. They formulated the problem as a Markov Decision Process (MDP) and developed a multi-agent DRL framework, i.e., SecBoost, using dueling double DQN and prioritized experience replay to maximize long-term secrecy energy efficiency. In [23], a joint task offloading and resource management problem was formulated to improve computation efficiency in MEC-enabled HetNets. The authors addressed the resulting MINLP using an Advantage Actor-Critic (A2C)-based algorithm (A2C-JTRA), which jointly optimizes offloading decisions, transmit power, and CPU frequency for MEC servers and end devices. In [24], a secrecy-oriented joint optimization of beamforming, artificial noise, and power control was proposed for Tera Hertz (THz)-enabled femtocells. The authors used a Multi-Agent Reinforcement Learning (MARL) framework and introduced two DRL-based methods, i.e., Deep Deterministic Policy Gradient-based Secrecy Maximization (DDPG-SM) and Asynchronous-Advantage-Actor-Critic based Secrecy Maximization (A3C-SM), to operate in continuous action spaces and enhance long-term secrecy performance. While these methods demonstrate promising results, they often struggle to generalize across dynamic user behaviors and high-dimensional state-action spaces.
II-B DRL for Networks
@@
DRL has been widely applied to network resource allocation tasks, where diverse user demands and time-varying conditions pose significant challenges to conventional optimization techniques [25]. In [26], the authors proposed a DRL-based Out-of-Order (OFO)-Friendly Multipath Scheduling (DOFMS) framework to ensure high-bandwidth and low-OFO transmission in mobile HetNets. The framework introduced a novel OFO metric for accurate evaluation, utilized a Double Deep Q-Network (DDQN) for dynamic scheduling, and incorporated an asynchronous learning module to handle variable path delays. In [27], a power control problem was formulated for spectrum-sharing HetNets to improve Global Energy Efficiency (GEE). The authors proposed a DRL-based cloud-edge collaborative framework that allows base stations to independently adjust transmit power, avoiding the scalability limitations of centralized optimization. In [28], a transformer-enhanced distributed DRL scheduling method was developed for heterogeneous IoT applications across edge and cloud computing environments. The method followed an actor-critic architecture, scaled across multiple distributed servers, and incorporated an off-policy correction mechanism to improve training stability. Despite these advances, existing DRL approaches often suffer from poor generalization and slow convergence in environments with heterogeneous user behaviors and high-dimensional state spaces.
II-C Transformer for Wireless Networks
@@
Transformer-based models have emerged as a promising solution for complex wireless networks, offering strong reasoning capabilities and adaptability for high-dimensional decision-making tasks. In [29], the authors proposed a Decision Transformer (DT)-based architecture for wireless resource management, aiming to improve sample efficiency and generalization over conventional DRL. By pre-training in the cloud and fine-tuning at the edge, the DT framework achieved faster convergence and better performance in intelligent reflecting surface-aided and unmanned aerial vehicle-assisted MEC scenarios. In [30], the authors studied deep learning-based wireless interference recognition under non-cooperative electromagnetic attacks. They proposed a multi-modal feature extraction framework with a dual transformer module and adaptive gradient modulation to improve recognition accuracy, convergence speed, and computational efficiency. In [31], a transformer-based model was introduced for Channel State Information (CSI) prediction in high-mobility scenarios. Using multi-head attention and positional encoding, the model effectively captured global channel features and outperformed traditional deep learning models in accuracy with reasonable complexity. Transformers have also been used to approximate value functions in DRL, including Q-networks and critic networks. For example, the authors in [32] proposed the Deep Transformer Q-Network (DTQN), which employs transformer decoders to encode agent observation histories, yielding more stable and efficient Q-value estimation in partially observable environments. Similarly, the authors in [33] introduced a transformer-based Soft Actor-Critic (SAC) variant that uses transformers to process $N$-step returns for improved value estimation in complex tasks. The Transformer Off-Policy Episodic RL (TOP-ERL) method [34] similarly leverages transformers to estimate value functions from segmented action sequences, improving the handling of long-term dependencies. Other related works explore planning-inspired architectures. Value Iteration Networks (VIN) [35] embed differentiable planning procedures into neural networks, while the Predictron architecture [36] conducts internal rollouts to estimate multi-step returns. TreeQN and ATreeC [37] incorporate tree search mechanisms into value networks to enable forward-looking reasoning. While these approaches focus primarily on temporal modeling or multi-step return estimation, ReaCritic takes a different direction. It introduces a transformer-based critic specifically designed to enhance bidirectional reasoning within single-step value estimation. This design enables more expressive and accurate critic outputs in complex, dynamic environments without relying on full rollout-based return prediction.
III System Model
@@ III-A Wireless Environment
[…]
where $m\in\{1,\ldots,M\}$, $\Delta_{x_{m}}^{\left(t\right)}\sim\mathcal{N}\left(0,\sigma^{2}\right)$ and $x_{m}^{\left(t\right)}$ is the normalized position of the $m_{\rm th}$ user at $t_{\rm th}$ step, and $\sigma$ controls the mobility intensity, allowing small random variations while ensuring that user positions remain within a bounded spatial domain. Each user’s transmission distance $D_{m}$ is derived from its normalized position and used to compute the large-scale path loss, modeled as
@@
Based on the SINR expressions, we derive three key performance metrics for each users in the HetNet: uplink and downlink data rates, energy efficiency, and service latency.
III-A1 Data Rate

### v2

# ReaCritic: Reasoning Transformer-based DRL Critic-model Scaling For Wireless Networks

@@ Abstract
Heterogeneous Networks (HetNets) pose critical challenges for intelligent management due to the diverse user requirements and time-varying wireless conditions. These factors introduce significant decision complexity, which limits the adaptability of existing Deep Reinforcement Learning (DRL) methods. In many DRL algorithms, especially those involving value-based or actor-critic structures, the critic component plays a key role in guiding policy learning by estimating value functions. However, conventional critic models often use shallow architectures that map observations directly to scalar estimates, limiting their ability to handle multi-task complexity. In contrast, recent progress in inference-time scaling of Large Language Models (LLMs) has shown that generating intermediate reasoning steps can significantly improve decision quality. Motivated by this, we propose ReaCritic, a reasoning transformer-based critic-model scaling scheme that brings reasoning-like ability into DRL. ReaCritic performs horizontal reasoning over parallel state-action inputs and vertical reasoning through deep transformer stacks. It is compatible with a broad range of value-based and actor-critic DRL algorithms and enhances generalization in dynamic wireless environments. Extensive experiments demonstrate that ReaCritic improves convergence speed and final performance across various HetNet settings and standard OpenAI Gym control tasks. The code of ReaCritic is available at https://github.com/NICE-HKU/ReaCritic.
@@
1 Introduction
The rapid growth of intelligent wireless services such as the Internet of Things (IoT), Mobile Edge Computing (MEC), and smart city applications is driving the evolution of future wireless networks toward large-scale Heterogeneous Networks (HetNets) [41]. These HetNets involve a large number of mobile users and dynamic environments with heterogeneous communication and computing resources [30]. The efficient management of these infrastructures faces three primary technical bottlenecks that hinder traditional resource management. First, the massive scale of mobile users and multi-tier base stations results in an explosive growth of state-action dimensionality whereby the search for optimal configurations becomes computationally prohibitive for conventional optimization techniques [15]. Second, the coexistence of heterogeneous network tiers creates intricate spatial-temporal couplings in which resource allocation at a specific access point exerts complex non-linear interference on neighboring nodes [27], thereby complicating the derivation of a global optimum in multi-user interference regimes. Third, the intrinsic non-stationarity of the wireless medium, driven by rapid user mobility, stochastic channel fading, and unpredictable traffic bursts, necessitates a decision-making framework capable of robust real-time adaptation rather than relying on static or predefined models [28].
As a result, HetNets must handle massive data traffic, ensure low-latency responses, and maintain energy efficiency under diverse service demands, which introduces fundamental challenges for scalable resource management and stable decision-making [1]. Specifically, the multifaceted requirements for diverse Quality-of-Service (QoS) levels, combined with the time-varying nature of wireless conditions, result in high-dimensional state spaces and complex user-resource interactions. Such properties fundamentally limit the scalability and generalization of traditional optimization and supervised learning-based methods, as they often fail to capture the latent dependencies required to maintain performance in non-stationary environments [54].
@@ 1 Introduction
Faced with these challenges, Deep Reinforcement Learning (DRL) has gained attention as a data-driven approach for autonomous network management in complex wireless environments [37]. By interacting with the environment and optimizing long-term rewards, DRL enables agents to adapt to dynamic network conditions without relying on predefined models [2]. Its success in domains such as intelligent control, robotics, and wireless resource management demonstrates its potential for handling large-scale, uncertain systems [9, 46, 50]. However, when applied to large-scale HetNets, DRL still faces key limitations not only due to environmental complexity but also due to internal architectural bottlenecks. Although DRL algorithms are capable of learning adaptive behaviors through trial-and-error interaction, many existing designs rely on simple Multi-Layer Perceptron (MLP)-based model structures that treat input states as flat feature vectors and apply direct approximations for policy or value functions [42]. This formulation fundamentally lacks the capacity to capture latent dependencies among state variables or to represent tradeoffs inherent in multi-objective decision-making. In heterogeneous wireless environments, where metrics such as latency, throughput, and energy efficiency are often tightly coupled and evolve asynchronously, such limitations can lead to unstable training, slow convergence, and poor performance in dynamic or unseen scenarios [20, 18]. Taking actor-critic-based DRL algorithms as an example, as shown in Fig. 1, the critic network estimates Q-values based on environmental states and the actions selected by the actor network.However, the high dimensionality and structural complexity of HetNet states and the action space make this mapping difficult and inaccurate, limiting actor updates and ultimately degrading overall DRL performance.
@@
Therefore, it is essential to enhance the scalability of DRL frameworks to better accommodate high-dimensional state spaces and dynamic network behaviors. One common strategy is to increase model depth or width, with the goal of improving the DRL’s capacity to interpret complex input features. However, scaling up the MLP parameter size alone often leads to overfitting, unstable convergence, and limited robustness in non-stationary environments. A more effective direction is to improve the DRL model through architectural adjustments that enable it to process structured observations and capture latent dependencies within the environment, beyond what can be achieved through conventional model size expansion. In this context, the architectural principles of Large Language Models (LLMs) offer a compelling source of inspiration. Recent advances in reasoning-capable LLMs have demonstrated strong capabilities in pattern abstraction and generalization across a wide range of tasks [17]. These LLMs are effective at modeling intricate relationships and long-term dependencies in structured or semi-structured data, opening opportunities for improving representation learning and decision-making in complex environments.
@@
  • Q1 (How): How can LLMs’ reasoning capability be effectively adapted to DRL settings without directly using pre-trained LLMs as external modules?
@@
To address Q1 (How), we adopt a Transformer-based reasoning architecture as a more stable and scalable alternative to conventional designs. Unlike traditional models, ReaCritic addresses HetNet challenges through two synergistic reasoning dimensions. First, the Horizontal Reasoning (HRea) module generates parallel reasoning tokens to explore a broader range of state-action correlations, effectively managing the breadth of high-dimensional input features. Second, the Vertical Reasoning (VRea) module utilizes stacked abstraction layers to perform hierarchical logic distillation, providing the depth needed to achieve stable convergence in non-stationary environments. To address Q2 (Where), we embed this reasoning-like capability into the critic module of the DRL framework. As the critic is responsible for value estimation and provides gradient signals to guide policy updates, it is the natural component for incorporating structured inference. By systematically scaling these reasoning dimensions, ReaCritic enables the critic to adapt its computational effort to the complexity of the task, thereby ensuring superior performance and robustness in large-scale wireless communication scenarios. This choice preserves the original policy structure while substantially improving the agent’s ability to model long-term dependencies, reason over complex state-action trajectories, and adapt to heterogeneous and time-varying network environments. Our proposed design, referred to as ReaCritic, is illustrated in Part C of Fig. 1. The algorithm can be centrally deployed, which avoids excessive computational load on edge devices and enables scalable coordination across heterogeneous users. In highly coupled environments, such centralized operation ensures globally consistent decisions and remains tractable for large, dynamic networks where performance is not heavily constrained by real-time latency.
@@
  • We propose ReaCritic, a reasoning transformer-based critic-model scaling architecture for DRL, inspired by inference-time scaling in LLMs. ReaCritic enables horizontal reasoning across parallel state-action token sequences and vertical reasoning through stacked transformer blocks, supporting adaptable Q value estimation in complex environments.
  • We conduct extensive experiments in a large-scale multi-user HetNet environment to evaluate ReaCritic. Compared with MLP-based critics and standard DRL baselines, ReaCritic achieves faster convergence, improved policy stability, and stronger adaptability under dynamic and wireless network conditions.
The remainder of this paper is organized as follows: Section 2 discusses related work in HetNets, DRL for networks, and transformer for wireless networks. Section 3 presents the wireless network system model, and gives the problem formulation of the optimization problem in the considered HetNet. Section 4 details the design of ReaCritic and the reasoning transformer-based critic-model scaling. Section 5 presents comprehensive experimental results and analysis. Finally, Section 6 concludes the paper with a summary of our key findings. A list of mathematical symbols frequently used in this paper is shown in Table I.
@@ 2.1 Heterogeneous Networks
HetNets have been studied using a range of techniques, including optimization [48], game theory [14], heuristics [26], and more recently, DRL [24]. These approaches aim to address the diverse user demands and dynamic resource constraints inherent in HetNets. Despite progress, the high dimensionality and coupling among decision variables make HetNets a particularly challenging yet critical domain for intelligent network management [39]. In [33], the authors proposed a secrecy-aware, energy-efficient scheme for HetNets, jointly optimizing power control, channel allocation, and beamforming under secrecy and Signal-to-Interference-plus-Noise Ratio (SINR) constraints. They formulated the problem as a Markov Decision Process (MDP) and developed a multi-agent DRL framework, i.e., SecBoost, using dueling Double Deep Q-Network (DDQN) and prioritized experience replay to maximize long-term secrecy energy efficiency. In [7], a joint task offloading and resource management problem was formulated to improve computation efficiency in MEC-enabled HetNets. The authors addressed the resulting MINLP using an Advantage Actor-Critic (A2C)-based algorithm (A2C-JTRA), which jointly optimizes offloading decisions, transmit power, and CPU frequency for MEC servers and end devices. In [32], a secrecy-oriented joint optimization of beamforming, artificial noise, and power control was proposed for Tera Hertz (THz)-enabled femtocells. The authors used a Multi-Agent Reinforcement Learning (MARL) framework and introduced two DRL-based methods, i.e., Deep Deterministic Policy Gradient-based Secrecy Maximization (DDPG-SM) and Asynchronous-Advantage-Actor-Critic based Secrecy Maximization (A3C-SM), to operate in continuous action spaces and enhance long-term secrecy performance. While these methods demonstrate promising results, they often struggle to generalize across dynamic user behaviors and high-dimensional state-action spaces.
2.2 DRL for Networks
@@
DRL has been widely applied to network resource allocation tasks, where diverse user demands and time-varying conditions pose significant challenges to conventional optimization techniques [34]. In [49], the authors proposed a DRL-based Out-of-Order (OFO)-Friendly Multipath Scheduling (DOFMS) framework to ensure high-bandwidth and low-OFO transmission in mobile HetNets. The framework introduced a novel OFO metric for accurate evaluation, utilized a Double DDQN for dynamic scheduling, and incorporated an asynchronous learning module to handle variable path delays. In [53], a power control problem was formulated for spectrum-sharing HetNets to improve Global Energy Efficiency (GEE). The authors proposed a DRL-based cloud-edge collaborative framework that allows base stations to independently adjust transmit power, avoiding the scalability limitations of centralized optimization. In [45], the authors proposed a transformer-enhanced distributed DRL scheduling method for heterogeneous IoT applications across edge and cloud computing environments. The method followed an actor-critic architecture, scaled across multiple distributed servers, and incorporated an off-policy correction mechanism to improve training stability. Despite these advances, existing DRL approaches often suffer from poor generalization and slow convergence in environments with heterogeneous user behaviors and high-dimensional state spaces.
2.3 Transformer for Wireless Networks
@@
Transformer-based models have emerged as a promising solution for complex wireless networks. In [52], the authors proposed a Decision Transformer (DT)-based architecture for wireless resource management, aiming to improve sample efficiency and generalization over conventional DRL. By pre-training in the cloud and fine-tuning at the edge, the DT framework achieved faster convergence and better performance in intelligent reflecting surface-aided and unmanned aerial vehicle-assisted MEC scenarios. In [43], the authors studied deep learning-based wireless interference recognition under non-cooperative electromagnetic attacks. They proposed a multi-modal feature extraction framework with a dual transformer module and adaptive gradient modulation to improve recognition accuracy, convergence speed, and computational efficiency. In [55], a transformer-based model was introduced for Channel State Information (CSI) prediction in high-mobility scenarios. Using multi-head attention and positional encoding, the model effectively captured global channel features and outperformed traditional deep learning models in accuracy with reasonable complexity. Transformers have also been used to approximate value functions in DRL, including Q-networks and critic networks. For example, the authors in [10] proposed the Deep Transformer Q-Network (DTQN), which employs transformer decoders to encode agent observation histories, yielding more stable and efficient Q-value estimation in partially observable environments. Similarly, the authors in [40] introduced a transformer-based Soft Actor-Critic (SAC) variant that uses transformers to process $N$-step returns for improved value estimation in complex tasks. The Transformer Off-Policy Episodic RL (TOP-ERL) method [22] similarly leverages transformers to estimate value functions from segmented action sequences, improving the handling of long-term dependencies. Other related works explore planning-inspired architectures. Value Iteration Networks (VIN) [38] embed differentiable planning procedures into neural networks, while the Predictron architecture [35] conducts internal rollouts to estimate multi-step returns. TreeQN and ATreeC [11] incorporate tree search mechanisms into value networks to enable forward-looking reasoning. While these approaches focus primarily on temporal modeling or multi-step return estimation, ReaCritic takes a different direction. It introduces a transformer-based critic specifically designed to enhance bidirectional reasoning within single-step value estimation. This design enables more expressive and accurate critic outputs in complex, dynamic environments without relying on full rollout-based return prediction.
3 System Model
@@ 3.1 Wireless Environment
$x_{m}^{(t+1)}=x_{m}^{(t)}+\Delta x_{m}^{(t)},\quad\Delta_{x_{m}}^{\left(t\right)}\sim\mathcal{N}\left(0,\sigma^{2}\right),$ (1)
where $m\in\{1,\ldots,M\}$ and $x_{m}^{\left(t\right)}$ is the normalized position of the $m_{\rm th}$ user at $t_{\rm th}$ step. This baseline mobility setting captures stochastic user movements and has been widely used for algorithm evaluation in wireless resource management studies [31]. The parameter $\sigma$ controls the mobility intensity, allowing small random variations within the spatial domain. Each user’s transmission distance $D_{m}$ is derived from its normalized position and used to compute the large-scale path loss, modeled as
@@
Based on the SINR expressions, we derive three key performance metrics for each user in the HetNet: uplink and downlink data rates, energy efficiency, and service latency.
3.1.1 Data Rate